Menu

University Libraries

All systems are online.

Minor Issue for Databases Unresolved

[2/9 @ 4:21 pm - Jeff]

Sanborn Maps is giving us some very slow response times. This is an ongoing issue with this product, but seems particularly bad today. I've opened a ticket with ProQuest.

This Minor Issue was reported on 2/9/16.

Update for Catalog Resolved

[2/9 @ 2:37 pm - Sara]

February Serial Solutions MARC Records 02/09/2016: We added 316 new serial solutions journal Marc records today. We now have 61,132 total in our collection.†​

This Update happened on 2/9/16.

Minor Issue for Homepage Unresolved

[2/8 @ 9:11 am - Matthew]

Banner is down on campus until at least 3pm today (Monday, 2/8), due to a DOS (Denial of Service) attack over the weekend. All systems that rely on connecting to Banner will be unavailable until Banner service is restored by IT. For us, this is limited to ScheduleIt, the software we use to book Knowledge Market appointments. I'll update once the services are restored.

This Minor Issue was reported on 2/8/16.

Outage for Databases Resolved

[2/5 @ 12:53 pm - Sara]

Mometrix has cut our access. We are working with them to get it restored. Until then, patrons will not have access.

[2/9 @ 2:32 pm - Sara]

access was restored the next day

This Outage was reported on 2/5/16 and resolved on 2/9/16.

Update for Catalog Resolved

[2/4 @ 8:50 am - Ben]

We deleted 1,322 EBL MARC records today.

This Update happened on 2/4/16.

Maintenance for Course Reserve Maintenance

[2/5 @ 11:59 pm - Mary]

Beginning on Friday, February 5, 2016 from 11:59PM PM EDT to 4:00 AM EDT Atlas Systems will be performing maintenance on the Ares (Course Reserves) hosted server's corresponding database server.

Web access to the Ares hosted server and client access to the Ares hosted serverís database may be unavailable. There will be no changes to the web pages or the data in the database.

This Maintenance is scheduled from 11:59pm 2/5/16 until 04:00am 2/6/16.

Update for Catalog Resolved

[2/2 @ 1:02 pm - Ben]

We added 490 Springer MARC records today.

This Update happened on 2/2/16.

Update for Catalog Resolved

[2/2 @ 12:20 pm - Ben]

We added 112 new NBER MARC records today.

This Update happened on 2/2/16.

Update for Catalog Resolved

[2/2 @ 10:54 am - Ben]

We added 2 new eBrary MARC records today and deleted 1,174.

This Update happened on 2/2/16.

Update for Catalog Resolved

[2/2 @ 10:12 am - Ben]

We deleted 92 Films on Demand MARC records today.

This Update happened on 2/2/16.